I have an issue with my Dr Who 5-opto target bank (that is not quite addressed by the many other opto-bank posts on Pinside).
I had a problem with switch 74 not working since I got the machine. After disassembly and a quick visual inspection, the opto receiver for that switch seemed to be damaged. I soldered in a replacement receiver, but it still does not register hits to the mushroom button. And ... now switch 73 stopped working as well.
However, when I test with the opto bank plugged into its connector (but not reinstalled in the MPF), I noticed if I jam a fingertip into the space between emitter & receiver, I can get both switches 73 & 74 to register (on the Switch Edges test)! So it appears that both emitter and receivers are working, but somehow even with the mushroom button stems blocking the IR beam, still enough light is reaching the receiver that it does not trigger.
Is that possible? Any suggested fixes?
I have checked the emitters with a digital camera--all seem to be shining brightly.
I have installed Swink's aftermarket channel stiffener (https://www.shapeways.com/product/W3GJK4YJQ/doctor-who-mushroom-target-channel-add-on) which has light baffles built-in, so light leak from an adjacent opto seems unlikely (and these are 2 of the middle targets, so light is not leaking in from the end of the channel. And my original construction paper light shield is intact.
I have a multimeter, but am not sure exactly how to test the receivers for proper resistance, etc. with the bank unplugged from machine.
Any suggestions for next steps? pinball_faz have you seen this before? Thanks!